OpenStudy (anonymous):

[y^2-2y/y^2+7y-18] X [y^2-81/y^2-11y+18] Simplify the rational expression.

OpenStudy (anikay):

\[\frac{ y^2-2y }{ y^2+7y-18 } * \frac{ y^2-8 }{ y^2-11y+18 }\]

OpenStudy (anikay):

is that correct?

OpenStudy (anonymous):

y^2-81. :P Now y2-2y = y(y-2), y2-81 = (y+9)(y-9), ok?

OpenStudy (anonymous):

you have to factor, multiply, and cancel out terms. I can't figure out how to factor and multiply

OpenStudy (anikay):

I can help but is my equation correct?

OpenStudy (anonymous):

yeah except on the top part of the second one it's an 81 not an 8

OpenStudy (anikay):

okay

OpenStudy (anikay):

okay \[\frac{ y(y-2) }{ (y-2)(y+9) } * \frac{ (y-9)^2 }{ (y-9)(y-2) }\]

OpenStudy (anikay):

do you understand how I got that?

OpenStudy (anonymous):

yes. thanks! so now do I just cancel?

OpenStudy (anikay):

yes

OpenStudy (anikay):

you should get \[\frac{ y }{ y+9 } * \frac{ y-9 }{ y-2 }\]

OpenStudy (anonymous):

okay, thanks. are there more steps?

OpenStudy (anikay):

if you multiply

OpenStudy (anikay):

(y^2-9y)/(y^2+7y-18)

OpenStudy (anonymous):

\(y^2 - 81 \neq (y - 9)^2\) \(y^2 - 81 = (y - 9)(y + 9)\)

OpenStudy (ryan123345):

Hey there! I see that you are new to OpenStudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! If you have any additional questions involving the website feel free to ask the people with the purples A's next to their name. They are the ambassadors and will answer all your questions regarding the site. And also read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Have a great day and I hope you have fun!

OpenStudy (anonymous):

\[\frac{ y^2-2y }{ y^2+7y-18 } * \frac{ y^2-81}{ y^2-11y+18 }\] \[\frac{ y(y - 2) }{ (y + 9)(y-2) } * \frac{ (y + 9)(y - 9)}{(y - 9)(y - 2)}\] \[\frac{y}{y + 9} * \frac{y + 9}{y - 2} = \frac{y}{y-2}\]

OpenStudy (anonymous):

I think you are supposed to divide out common factors to cancel. like y-2 and y-2

OpenStudy (anonymous):

\[\frac{ y(y - 2) }{ (y + 9)(y-2) } * \frac{ (y + 9)(y - 9)}{(y - 9)(y - 2)}=\] \[\frac{ y(y - 2)(y + 9)(y - 9) }{ (y + 9)(y-2)(y - 9)(y - 2) }\] Canceling stuff \[\frac{ y\cancel{(y - 2)}\cancel{(y + 9)}\cancel{(y - 9)}}{ \cancel{(y + 9)}\cancel{(y-2)}\cancel{(y - 9)}(y - 2) }=\frac{y}{y - 2}\]